# Ferromagnetic Resonance in a Magnetically Dilute Percolating Ferromagnet: An Experimental and Theoretical Study

### v2

@@ III.4 FMR linewidth and relaxation in (Ga,Mn)N
The investigation of the FMR linewidth ($\Delta H$) can give valuable information on the relaxation rate of the magnetization as well as potential inhomogeneities of the studied system. $\Delta H$ contains the contributions originating from the intrinsic damping of the magnetization precession, $\Delta H_{homo}$, and the contribution related to magnetic and/or structural inhomogeneities in the system. In the $(1\bar{1}00)$ rotation plane $\Delta H_{homo}$ simplifies to [81]:
@@
The fit of experimental data with the sum of Eqs.(8) and (9) presented by the solid line in Fig. 6 gives $\Delta H_{homo}=2070$ Oe and $\Delta H_{2}=63$ Oe. In the fit the angular dependence of $\Delta H_{homo}$ is neglected since the maximum deviation of $\phi$ from $\phi_{H}$ (at $\phi_{H}=50$ deg) leads only to 1.5% line broadening, which falls within the experimental error. It is evident that the main contribution to the linewidth is due to intrinsic damping. The Gilbert parameter determined from Eq. 8 is about 4×$\times 10^{8}\leavevmode\nobreak\ s^{-1}$ for H along [0001] at 12 K and does not differ much from the values found for typical ferromagnetic metals [38]. The reason for the untypically broad linewidth is the low value of magnetization [$M\simeq 45$ emu/cm3, as indicated in Fig. 2 (b)], which enters in the denominator of Eq.(8). This explains also the unusually high value of the parameter $\alpha_{G}=\frac{\hbar G}{g\mu_{B}M}$ = 0.68 as compared to typical values observed in diluted magnetic systems and metals [38, 83, 84].
@@ IV Summary
In this study, we have thoroughly investigated the magnetic anisotropy and relaxation dynamics in percolating ferromagnetic semiconductor Ga1-xMnxN film with $x\sim 8$% using ferromagnetic resonance (FMR) and magnetometry. Our results confirm the ferromagnetic nature of the material, dominated by Mn3+ ions, which for this Mn dilution yields a clear ferromagnetic response below $T_{\mathrm{C}}=12$ K.
While the FMR signal is absent at low temperatures ($T<9$ K) due to excessive line broadening that prevents detection, the signal persists up to 70 K, revealing the presence of non-percolating FM clusters (macrospins) above $T_{\mathrm{C}}$. The temperature dependence of the signal intensity mirrors that of the magnetization up to 55 K, suggesting that the magnetic characteristics of the clusters remain relatively stable across this temperature range. Furthermore, the evolution of the FMR signal with temperature provides direct evidence of the progressive fragmentation of the percolating ferromagnetic network into isolated macrospins.
A crucial observation from our study is that while (Ga,Mn)N exhibits percolative ferromagnetism, it retains essential features of conventional ferromagnetic materials, particularly in its resonance behavior. The dominance of uniaxial anisotropy with the easy axis perpendicular to the $c$-axis across the whole temperature range studied here confirms the dominant role of Mn3+ single ion magnetic anisotropy. Despite prior reports suggesting the possible influence of static Jahn-Teller distortions, our FMR results do not show evidence of such effects, implying a transition from a static to a dynamic Jahn-Teller regime at the temperatures studied here.
Our atomistic spin model simulations qualitatively reproduce the experimentally observed trends, providing further insights into the nature of ferromagnetic interactions in this dilute system. While the agreement between experiment and simulation is satisfactory, further refinements in modeling Mn-Mn interactions, especially in the presence of short-range disorder, are necessary for more accurate predictions. The combined use of SQUID and FMR techniques allows us to bridge the gap between static and dynamic magnetic properties, illustrating the complex yet highly informative nature of percolating ferromagnetic semiconductors.
In summary, our study confirms that FMR serves as a powerful complementary technique to magnetometry for probing the magnetic behavior of percolating ferromagnetic systems like (Ga,Mn)N. The persistence of FMR signals above $T_{\mathrm{C}}$ highlights the role of non percolating ferromagnetic clusters. Our findings provide new insights into the interplay between magnetization dynamics, anisotropy, and percolation effects in (Ga,Mn)N, contributing to the broader understanding of percolating ferromagnetic systems and their potential relevance for spintronic applications, even well above $T_{\mathrm{C}}$.
@@ Appendix B Simulation details
The magnetocrystalline energy is composed of (uniaxial) trigonal and Jahn-Teller anisotropies. The latter is approximated by a cubic form [51, 33], with anisotropy axes aligned along the three JT distortion axes, defined after Ref. 31: $e_{A}^{JT}=\left[\sqrt{\frac{2}{3}},0,\sqrt{\frac{1}{3}}\right]$, $e_{B}^{JT}=\left[-\sqrt{\frac{1}{6}},-\sqrt{\frac{1}{2}},\sqrt{\frac{1}{3}}\right]$, $e_{C}^{JT}=\left[-\sqrt{\frac{1}{6}},\sqrt{\frac{1}{2}},\sqrt{\frac{1}{3}}\right]$.
@@
Technically, in the sLLG simulations, the code is parallelized by dividing the simulation region into sections, with each processor handling a specific part of the system. Computations are accelerated using graphics processing units (GPUs), and Euler’s method is employed as the integration scheme after checking that it yields the same results as the Heun’s scheme and that all results are stable. Since a higher temperature introduces more disturbance in the system, the initialization and averaging steps of the sLLG simulations are adjusted as specified in table 1.
@@ 2.50 $2\times 10^{7}$ $6\times 10^{7}$
Appendix C Limitation of the model
While the atomistic spin model employed in this study captures the essential features of the magnetization behavior in (Ga,Mn)N, several simplifying assumptions were made, which limit the accuracy of the simulations—particularly in the low-field regime where the Zeeman energy is insufficient to overcome subtle anisotropy effects. The primary limitations of the current model are as follows:
  • Neglect of Mn–Mn pair anisotropy. Pairwise anisotropic interactions between neighboring Mn ions are not included in the present model. These interactions have been shown to influence magnetic behavior in dilute systems and are currently under intensive investigation in a follow-up study [32]. These effects may be important, as, assuming purely statistical Mn distribution, for $x=0.08$ as much as $1-(1-x)^{12}$ Mn ions, that is 63% of all ions must have at least one nearest-neighbor. Their inclusion is expected to improve the fidelity of the simulated remnant magnetization and coercivity.
  • Cubic approximation of Jahn–Teller anisotropy. In reality, the Jahn–Teller effect leads to a tetragonal distortion along one of the cubic ⟨100⟩ axes, with the trigonal ⟨111⟩ axis aligned with the crystallographic $\mathbf{c}$-axis. This implies that the distortion is active along only one of the local axes at a time ($e_{A}$, $e_{B}$, or $e_{C}$ defined in the previous section), as captured in direct quantum mechanical calculations and crystal field model [43, 31]. However, such directional specificity (lowering of the site symmetry) cannot be implemented within the classical sLLG framework. Therefore, we approximate the Jahn–Teller contribution using an effective cubic anisotropy term (the second part of Eq. 10), i.e., acting equally along the three crystallographic directions, in contrast to the Jahn–Teller distortion, which selects only one axis at a time.
  • Assumption of a perfectly random Mn distribution. The spatial distribution of Mn ions is assumed to be statistically random, without accounting for clustering or local concentration fluctuations that may occur during growth [76]. Experimental and theoretical studies suggest that such inhomogeneities can significantly modify the local anisotropy landscape and spin connectivity, particularly in the dilute regime [17, 18].

### v3

@@ III.4 FMR linewidth and relaxation in (Ga,Mn)N
The investigation of the FMR linewidth ($\Delta H$) can give valuable information on the relaxation rate of the magnetization as well as potential inhomogeneities of the studied system. $\Delta H$ contains the contributions originating from the intrinsic damping of the magnetization precession, $\Delta H_{homo}$, and the contribution related to magnetic and/or structural inhomogeneities in the system.
In the $(1\bar{1}00)$ rotation plane $\Delta H_{homo}$ simplifies to [77]:
@@
The fit of experimental data with the sum of Eqs.(8) and (9) presented by the solid line in Fig. 6 gives $\Delta H_{homo}=2070$ Oe and $\Delta H_{2}=63$ Oe. In the fit the angular dependence of $\Delta H_{homo}$ is neglected since the maximum deviation of $\varphi$ from $\varphi_{H}$ (at $\varphi_{H}=50$ deg) leads only to 1.5% line broadening, which falls within the experimental error. It is evident that the main contribution to the linewidth is due to intrinsic damping. The Gilbert parameter determined from Eq. 8 is about 4×$\times 10^{8}\leavevmode\nobreak\ s^{-1}$ for H along [0001] at 12 K and does not differ much from the values found for typical ferromagnetic metals [34]. The reason for the untypically broad linewidth is the low value of magnetization [$M\simeq 45$ emu/cm3, as indicated in Fig. 2 (b)], which enters in the denominator of Eq.(8). This explains also the unusually high value of the parameter $\alpha_{G}=\frac{\hbar G}{g\mu_{B}M}$ = 0.68 as compared to typical values observed in diluted magnetic systems and metals [34, 79, 80].
@@ IV Summary
In this study, we have investigated the magnetic anisotropy and relaxation dynamics in percolating ferromagnetic semiconductor Ga1-xMnxN with $x\sim 8$% using ferromagnetic resonance (FMR) and SQUID magnetometry. Our results confirm the ferromagnetic nature of the material, dominated by Mn3+ ions, with a well-defined transition at $T_{\mathrm{C}}=12$ K. Although the FMR signal vanishes below $T\approx 9$ K due to excessive line broadening, it persists up to 70 K. This persistence, alongside the $M(T)$ behavior, indicates the presence of stable, non-percolating ferromagnetic clusters (macrospins) well above $T_{\mathrm{C}}$, and supports a scenario of progressive fragmentation of the percolating FM network with increasing temperature.
While exhibiting percolative ferromagnetism (Ga,Mn)N retains key features of conventional ferromagnets, including clear resonance behavior and dominant uniaxial magnetic anisotropy with an easy axis perpendicular to the $c$-axis, attributed to Mn3+ single-ion anisotropy. No signatures of static Jahn–Teller distortions are observed, implying a transition to a dynamic regime in the studied temperature range. Atomistic spin simulations qualitatively reproduce the experimental results and offer insights into the nature of Mn–Mn interactions in this dilute system. While the model captures the main trends, further refinements are required to fully account for local disorder and low-field discrepancies.
Overall, our results confirm that FMR serves as a powerful complementary technique to magnetometry in studying percolating ferromagnetic systems. The coexistence of non percolating ferromagnetic clusters above $T_{\mathrm{C}}$ provides a deeper understanding of the complex interplay between magnetization dynamics, anisotropy, and percolation effects, with potential implications for spintronic applications even well above the Curie temperature.
@@ Appendix B Simulation details
The magnetocrystalline energy is composed of (uniaxial) trigonal and Jahn-Teller anisotropies. The latter is approximated by a cubic form [44, 30], with anisotropy axes aligned along the three JT distortion axes, defined after Ref. 28: $e_{A}^{JT}=\left[\sqrt{\frac{2}{3}},0,\sqrt{\frac{1}{3}}\right]$, $e_{B}^{JT}=\left[-\sqrt{\frac{1}{6}},-\sqrt{\frac{1}{2}},\sqrt{\frac{1}{3}}\right]$, $e_{C}^{JT}=\left[-\sqrt{\frac{1}{6}},\sqrt{\frac{1}{2}},\sqrt{\frac{1}{3}}\right]$.
@@
Technically, in the sLLG simulations, the code is parallelized by dividing the simulation region into sections, with each processor handling a specific part of the system. Computations are accelerated using graphics processing units (GPUs), and Euler’s method is employed as the integration scheme after checking that it yields the same results as the Heun’s scheme (see Appendix C), and that all results are stable. Since a higher temperature introduces more disturbance in the system, the initialization and averaging steps of the sLLG simulations are adjusted as specified in table 1.
@@ 2.50 $2\times 10^{7}$ $6\times 10^{7}$
Appendix C Integration scheme validation
In the sLLG simulations the Euler integration scheme has been applied. To ensure that our choice does not affect the reliability of the modeling results, we perform a comparative analysis with the Heun method. Simulations are conducted using identical parameters and timestep $\Delta t=5\cdot 10^{-6}$ ns for both schemes. Fig. 9 shows a side-by-side comparison of magnetization curves and FMR spectra obtained by both methods.
Figure 9: (Color online) Comparison between Euler (open rectangles) and Heun (full circles) integration schemes used during stochastic Landau–Lifshitz–Gilbert simulations for: (a) magnetization curves $M(H)$ at $T=5$ K for $\mathbf{H}\perp\mathbf{c}$, (b) $M(H)$ for $\mathbf{H}\parallel\mathbf{c}$, and (c) FMR spectrum at $T=4$ K with $\mathbf{H}\parallel\mathbf{c}$. The green diamonds in panels (a) and (b) mark the difference between the results of the two integration schemes, while the solid lines in panel (c) shows the results of the fit of the Eq. 13 to the numerically simulated points.
As illustrated, both the static magnetization and dynamic FMR results are in very good agreement for the two schemes. The relative differences in resonance field ($H_{R}$) and linewidth ($\Delta H$) remain within 5%.
Appendix D Limitation of the model
While the atomistic spin model employed in this study captures the essential features of the magnetization behavior in (Ga,Mn)N, several simplifying assumptions have been applied, which limit the accuracy of the simulations—particularly in the low-field regime where the Zeeman energy is insufficient to overcome subtle anisotropy effects. The primary limitations of the current model are as follows:
  • Neglect of Mn–Mn pair anisotropy: Pairwise anisotropic interactions between neighboring Mn ions are not included in the present model. These interactions have been shown to influence magnetic behavior in dilute systems and are currently under intensive investigation in a follow-up study [29]. These effects may be important, as, assuming purely statistical Mn distribution, for $x=0.08$ as much as $1-(1-x)^{12}$ Mn ions, that is 63% of all ions must have at least one nearest-neighbor. Their inclusion is expected to improve the fidelity of the simulated remnant magnetization and coercivity.
  • Cubic approximation of Jahn–Teller anisotropy: In reality, the Jahn–Teller effect leads to a tetragonal distortion along one of the cubic ⟨100⟩ axes, with the trigonal ⟨111⟩ axis aligned with the crystallographic $\mathbf{c}$-axis. This implies that the distortion is active along only one of the local axes at a time ($e_{A}$, $e_{B}$, or $e_{C}$ defined in the previous section), as captured in direct quantum mechanical calculations and crystal field model [36, 28]. However, such directional specificity (lowering of the site symmetry) cannot be implemented within the classical sLLG framework. Therefore, we approximate the Jahn–Teller contribution using an effective cubic anisotropy term (the second part of Eq. 10), i.e., acting equally along the three crystallographic directions, in contrast to the Jahn–Teller distortion, which selects only one axis at a time.
  • Assumption of a perfectly random Mn distribution: The spatial distribution of Mn ions is assumed to be statistically random, without accounting for clustering or local concentration fluctuations that may occur during growth [72]. Experimental and theoretical studies suggest that such inhomogeneities can significantly modify the local anisotropy landscape and spin connectivity, particularly in the dilute regime [15, 16].